# GP+: A Python Library for Kernel-based learning via Gaussian Processes

### v1

@@ 1 Introduction
Gaussian processes (GPs) are indispensable building blocks of many powerful probabilistic frameworks such as Bayesian optimization (BO) [1, 4, 5, 6, 7, 8, 9, 10, 11, 2, 3], function and operator learning [12, 13, 14, 15, 16, 17], data-driven calibration of expensive simulations [18, 19, 20, 21, 22, 23], and multi-fidelity (MF) modeling [24, 25, 26, 27, 28]. The recent software and hardware developments, combined with the new means of data collection and the society’s drive to tackle ever-challenging goals, have sparked significant innovations in the broad field of machine learning (ML). GPs have also substantially benefited from these advancements and recent works have enabled them to leverage GPUs [29, 30] and to accommodate high dimensions [31, 32, 33, 34, 35], large datasets [36, 37, 38, 39, 40, 41, 42, 43, 44], or disjoint feature spaces that have both qualitative and quantitative variables [45, 46, 47, 48, 16]. In line with these advancements, in this paper we introduce GP+ which is a modular and user-friendly library that aims to empower researchers and practitioners in harnessing the full potential of GPs for a wide range of applications such as probabilistic metamodeling (aka emulation), single- and multi-fidelity Bayesian optimization (SFBO and MFBO), kernel-based generalized MF modeling, inverse parameter estimation, anomaly detection, and sensitivity analyses.
@@
As reviewed in Section 3, there are quite a few existing libraries for GP modeling and in fact we leverage one of them (i.e., GPyTorch [30]) in developing ours. While these libraries have been successfully used in many applications across sciences and engineering, we believe a few distinct features set GP+ aside. Firstly, we design parametric covariance and prior mean functions that automatically enable GPs to fuse and emulate multi-source datasets, detect anomalies, find calibration parameters of computer models (i.e., inverse parameter estimation), and handle categorical variables. These functions leverage kernel-based nonlinear manifold learning, provide interpretable solutions (see Section 5 for multiple examples), and their parameters are all jointly learnt via the maximum a posteriori (MAP) method. Secondly, we provide a unified platform to use GPs for many tasks such as SFBO and MFBO, probabilistic regression, and sensitivity analysis. As shown in Section 5, all of these functionalities are achieved via a few lines of codes. As an example, we develop a hyperparameter estimation routine based on the method of continuation [49] which, at the expense of slightly higher computational costs, provides more numerical stability and accuracy. This method is accessible by merely changing the default optimization settings in GP+. Lastly, GP+ is accompanied with a rich set of datasets (from engineering applications) and benchmark analytic examples that can be used by researchers in evaluating the performance of emulation techniques beyond GPs.
@@ 2 Background on Gaussian Processes
where $\boldsymbol{\mu}^{*}_{q}=\boldsymbol{\mu}_{q}+\boldsymbol{\Sigma}_{qn}\boldsymbol{\Sigma}^{-1}_{nn}(\boldsymbol{y}_{n}-\boldsymbol{\mu}_{n})$ is the conditional mean vector and $\boldsymbol{\Sigma}^{*}_{qq}=\boldsymbol{\Sigma}_{qq}-\boldsymbol{\Sigma}_{qn}\boldsymbol{\Sigma}^{-1}_{nn}\boldsymbol{\Sigma}_{nq}$ is the conditional covariance matrix. Similarly, in interpolation with GPs one first assumes that the given data $\boldsymbol{y}_{n}$ and the to-be-predicted values $\boldsymbol{y}_{q}$ are jointly Gaussian and then infers the latter via Equation 2. However, the mean vectors and covariance matrices in Equation 1 are unknown and thus Equation 2 can be used for prediction only after we $(1)$ endow the underlying GP with a parametric mean function and a parametric covariance function (or kernel), and $(2)$ estimate the parameters of these two sets of functions [^2] In a fully Bayesian setting, instead of estimating the parameters, their posterior distributions are obtained and predictions on $\boldsymbol{y}_{q}$ require marginalization with respect to these distributions. Due to the significantly higher computational costs of fully Bayesian techniques and their marginal accuracy improvements in the case of GPs, we recommend and use MAP..
More formally, assume the training dataset $\mathopen{}\left\{\boldsymbol{x}^{(i)},y^{(i)}\right\}\mathclose{}_{i=1}^{n}$ is given where $\boldsymbol{x}=[x_{1},...,x_{dx}]^{T}\in\mathbb{X}\subset\mathbb{R}^{dx}$ and $y^{(i)}=y(\boldsymbol{x}^{(i)})\in\mathbb{R}$ denote the inputs and response [^3] We focus on regression problems whose output dimensionality is one but note that GPs can handle multi-response or multi-task problems as well [50, 51, 52]., respectively. Given $\boldsymbol{y}=[y^{(1)},\cdots,y^{(n)}]^{T}$ and $\boldsymbol{X}$ whose $i^{th}$ row is $(\boldsymbol{x}^{(i)})^{T}$, our goal is to predict $y(\boldsymbol{x}^{*})$ at the arbitrary point $\boldsymbol{x}^{*}\in\mathbb{X}$. Following the above description, we assume $\boldsymbol{y}=[y^{(1)},\cdots,y^{(n)}]^{T}$ is a realization of a GP with the following parametric mean and covariance functions:
@@
where $\omega_{i}\in\mathcal{R}$ [^4] To ensure numerical stability, $\omega_{i}$ is typically bounded to a subset of $\mathcal{R}$, e.g., $[-10,4]$., $p\in[1,2]$, $\nu\in\{\frac{1}{2},\frac{3}{2},\frac{5}{2}\}$, $K_{\nu}$ is the modified Bessel function of the second kind, and $\Gamma$ is the gamma function. The inductive bias that the kernels in Section 2 encode into the learning process is that close-by input vectors $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$ have similar (i.e., correlated) output values. The degree of this correlation is quantified by the interpretable length-scale (aka roughness) parameters where the magnitude of $10^{\omega_{i}}$ is directly related to the response fluctuations along $x_{i}$.
@@
where $c(\boldsymbol{x}^{*},\boldsymbol{X};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})$ is a $1\times n$ row vector with entries $c_{i}=c(\boldsymbol{x}^{*},\boldsymbol{x}^{(i)};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})$ and its transpose is $c(\boldsymbol{X},\boldsymbol{x}^{*};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})$. Equations 9a and 9b can be straightforwardly extended to predict the response distribution for a batch of samples denoted by $\boldsymbol{X}^{*}$:
@@
$\text{cov}\mathopen{}\left(\mathopen{}\left(y\boldsymbol{X}^{*}\right)\mathclose{},y(\boldsymbol{X}^{*})\right)\mathclose{}=c(\boldsymbol{X}^{*},\boldsymbol{X}^{*};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})-c(\boldsymbol{X}^{*},\boldsymbol{X};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})\boldsymbol{C}^{-1}c(\boldsymbol{X},\boldsymbol{X}^{*};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})$ (10b)
The above formulations build interpolating GPs. To handle datasets with noisy observations, the nugget or jitter parameter, denoted by $\delta$ [53, 12, 54], is used where $\boldsymbol{C}$ is replaced by $\boldsymbol{C}_{\delta}=\boldsymbol{C}+\delta\boldsymbol{I}_{nn}$ where $\boldsymbol{I}_{n\texttimes n}$ is the $n\times n$ identity matrix (with this adjustment, the stationary noise variance estimated by the GP is $\widehat{\delta}$) [^6] Some recent works [55, 56, 57] apply the nugget directly to $\boldsymbol{R}$ but herein we adhere to [53] and add $\delta$ to $\boldsymbol{C}$.. In addition to modeling stationary noise, the nugget parameter is also used to mitigate the numerical issues associated with $\boldsymbol{C}$. That is, even with noise-free $\boldsymbol{y}$, $\boldsymbol{C}_{\delta}$ is used while minimizing Equation 8 to ensure the correlation matrix is always invertible. When the nugget parameter is used for fitting a GP to noisy observations, Equation 1 takes on the following form:
@@
$\text{cov}\mathopen{}\left(\mathopen{}\left(y\boldsymbol{X}^{*}\right)\mathclose{},y(\boldsymbol{X}^{*})\right)\mathclose{}=c(\boldsymbol{X}^{*},\boldsymbol{X}^{*};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})-c(\boldsymbol{X}^{*},\boldsymbol{X};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})\boldsymbol{C}^{-1}_{\delta}c(\boldsymbol{X},\boldsymbol{X}^{*};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})+\widehat{\delta}\boldsymbol{I}.$ (12b)
We highlight that Equation 12b does not consider the additional uncertainties incurred by estimating the parameters of the mean and covariance functions (note that Equation 2 assumed the mean vector and covariance matrices are known). These additional uncertainties can be quantified by building and using the GP within a Bayesian framework where sampling methods (e.g., Markov Chain Monte Carlo or MCMC) are required for marginalization as closed form expressions are only available for specific cases (see [58] for an example). Since such sampling methods are typically expensive and the provided benefits are marginal, MAP is frequently employed in GP modeling.
@@ 3 Related Works
Many open-source GP libraries have been recently developed and in this section we review some of the most well-known ones. One of the earliest open-source GP packages is TreedGP [59] which is developed in R and primarily aims to address the stationarity and scalability issues of GPs. In particular, TreedGP recursively partitions the input space via parallel and axes-aligned boundaries [60] and then endows each partition with a GP whose covariance function is stationary. TreedGP uses Bayesian averaging to combine these GPs which is particularly important for obtaining smooth predictions on the partition boundaries. The major limitations of TreedGP are its non-differentiability on the boundaries, high computational costs (as the Bayesian analyses rely on MCMC [61]), reliance on trees which can only partition the input space with axis-aligned boundaries [62], and inability to efficiently handle categorical features in small-data applications.
GPfit [57] and GPM [56] are also R packages and they are primarily designed to improve the hyper-parameter optimization process at the expense of increased computational costs. GPfit has a multi-step pre-processing stage that aims to improve the quality of the initial points that are used via L-BFGS in minimizing Equation 8. Unlike GPfit, GPM develops a multi-step continuation-based strategy to increase both the robustness and accuracy of the optimization process. In particular, GPM indirectly controls $\delta$ via the auxiliary parameter $\epsilon$ that puts a lower bound on the smallest eigenvalue of the correlation matrix [^7] The rationale behind this choice is that the smallest eigenvalue of $\boldsymbol{R}$ can sometimes be negative due to numerical issues.. GPM first uses a large value for $\epsilon$ (e.g., $10^{-2}$) and minimizes Equation 8 while requiring the smallest eigenvalue of $\boldsymbol{R}$ to always be larger than the imposed $\epsilon$. In addition to guaranteeing numerical robustness, this requirement dramatically smooths the profile of the objective function and hence most (if not all) optimizations quickly converge to the same solution. Then, GPM relaxes the constraint on $\boldsymbol{R}$ (e.g., $\epsilon=10^{-3}$) and repeats the optimization while using the solution(s) of the previous step as the initial guess(es) in the current step, see Figure 1. These steps are continued until the minimum value of $\epsilon$ is reached and then the parameters of the final GP are chosen by identifying the step (or $\epsilon$) at which the leave-one-out cross-validation (LOO-CV) error of the model is minimized.
@@
GPy [63] is a popular object-oriented library that is implemented based on numeric Python (NumPy) by the Sheffield machine learning group. GPy provides a number of basic and advanced functionalities for GP regression that include multi-output learning and non-Gaussian likelihood functions which are accompanied with Laplace approximation [64] and expectation propagation since exact inference with non-Gaussian likelihoods is not tractable. However, GPy does not fully leverage modern hardware capabilities (e.g., GPU acceleration) and integration with deep neural networks (DNNs) which are increasingly crucial in contemporary GP applications. It also lacks some of the most recent advancements that enable GPs to accommodate high dimensions or categorical features.
One of the first open-source GP libraries that supports GPU acceleration and leverages automatic differentiation is GPflow [29] which is based on Tensorflow and has an object-oriented Python front-end. GPflow supports regression and classification problems, uses variationally sparse methods for scalability to large data, and provides both Bayesian and point-estimate-based inference classes for Gaussian and non-Gaussian likelihoods. While GPflow has significant capabilities, it lacks some of the key recent advancements in GPs such as natural integration with SFBO or MFBO frameworks, fusing multi-source data, calibrating unknown parameters, or directly supporting categorical variables.
Perhaps the most widely used open-source package for GP modeling is GPytorch [30] which accommodates a wealth of functionalities such as deep kernel [65] and multi-task/multi-output learning [50, 51, 52], dimensionality reduction via latent variable GPs [66, 67], and variational and approximate modeling for handling non-Gaussian likelihoods (e.g., in classification) or large datasets. A particular feature of GPytorch is its use of preconditioned conjugate gradients in accelerating the expensive computations (i.e., matrix inversion and determinant calculation) associated with exact GPs in Equations 8 and 10b. This feature is coined blackbox matrix-matrix multiplication (BBMM) and is uniquely suited for GPU-based computations. GPytorch relies on Pyro [68] for probabilistic programming, forms the backbone of BoTorch which is an open-source Python library for BO, and has specialized kernels for MF modeling or handling categorical variables. Given the versatility and robustness of GPytorch, we use some of its core functionalities to build GP+ while addressing the limitations of GPytorch in handling categorical features, directly solving inverse problems, learning from MF datasets, or interfacing with SFBO/MFBO engines.
Other notable open-source GP libraries that we mention in passing include those of Ambikasaran et al. [69] which hierarchically factors the covariance matrix [^8] As long as it is built with specific covariance functions such as the Gaussian or Matérn. into a product of block low-rank updates of the identity matrix to accelerate matrix inversion and determinant calculation (their method loses accuracy for $dx>3$), Vanhatalo et al. [70] which is a Matlab library that integrates various elementary computational tools (e.g., sparse approximation) for GP-based regression, and GPML [12] which is also a Matlab library and has been widely used in a wide range of applications.

### v2

@@ 1 Introduction
Gaussian processes (GPs) are indispensable building blocks of many powerful probabilistic frameworks such as Bayesian optimization (BO) [RN578, RN1242, RN1375, RN1392, RN1576, RN1577, RN1632, RN1761, RN1789, RN1794, RN1845], function and operator learning [RN332, RN1881, RN1886, RN1890, RN1559, RN1558], data-driven calibration of expensive simulations [RN271, RN272, RN334, RN337, RN705, RN841], and multi-fidelity (MF) modeling [RN1266, RN1275, RN1838, RN1786, RN791]. The recent software and hardware developments, combined with the new means of data collection and the society’s drive to tackle ever-challenging goals, have sparked significant innovations in the broad field of machine learning (ML). GPs have also substantially benefited from these advancements and recent works have enabled them to leverage GPUs [RN1894, RN1270] and to accommodate high dimensions [RN321, RN530, RN713, RN1228, RN1412], large datasets [RN718, RN893, RN503, RN698, RN818, RN1574, RN1479, RN1573, RN347], or disjoint feature spaces that have both qualitative and quantitative variables [RN1075, RN1215, RN1214, RN1492, RN1559]. In line with these advancements, in this paper we introduce GP+ which is a modular and user-friendly library that aims to empower researchers and practitioners in harnessing the full potential of GPs for a wide range of applications such as emulation (i.e., probabilistic metamodeling), single- and multi-fidelity Bayesian optimization (SFBO and MFBO), kernel-based generalized MF modeling, inverse parameter estimation, anomaly detection, and sensitivity analyses.
@@
As reviewed in Section 3, there are quite a few existing libraries for GP modeling and in fact we leverage one of them (i.e., GPyTorch [RN1270]) in developing ours. While these libraries have been successfully used in many applications across sciences and engineering, we believe a few distinct features set GP+ aside. First, we design parametric covariance and prior mean functions that automatically enable GPs to fuse and emulate multi-source datasets, detect anomalies, find calibration parameters of computer models (i.e., inverse parameter estimation), and handle categorical variables. These functions leverage kernel-based nonlinear manifold learning, provide interpretable solutions (see Section 5 for multiple examples), and their parameters are all jointly learnt via the maximum a posteriori (MAP) method. Second, we provide a unified platform to use GPs for many tasks such as SFBO and MFBO, probabilistic regression, and sensitivity analysis. As shown in Section 5, all of these functionalities are achieved via a few lines of codes. As an example, we develop a hyperparameter estimation routine based on the method of continuation [RN1062] which, at the expense of slightly higher computational costs, provides more numerical stability and accuracy. This method is accessible by merely changing the default optimization settings in GP+. Lastly, GP+ is accompanied with a rich set of datasets (from engineering applications) and benchmark analytic examples that can be used by researchers in evaluating the performance of emulation techniques beyond GPs.
@@ 2 Background on Gaussian Processes
where $\boldsymbol{\mu}^{*}_{q}=\boldsymbol{\mu}_{q}+\boldsymbol{\Sigma}_{qn}\boldsymbol{\Sigma}^{-1}_{nn}(\boldsymbol{y}_{n}-\boldsymbol{\mu}_{n})$ is the conditional mean vector and $\boldsymbol{\Sigma}^{*}_{qq}=\boldsymbol{\Sigma}_{qq}-\boldsymbol{\Sigma}_{qn}\boldsymbol{\Sigma}^{-1}_{nn}\boldsymbol{\Sigma}_{nq}$ is the conditional covariance matrix. Similarly, in interpolation with GPs one first assumes that the given data $\boldsymbol{y}_{n}$ and the to-be-pblackicted values $\boldsymbol{y}_{q}$ are jointly Gaussian and then infers the latter via Equation 2. However, the mean vectors and covariance matrices in Equation 1 are unknown and thus Equation 2 can be used for pblackiction only after we $(1)$ endow the underlying GP with a parametric mean function and a parametric covariance function (or kernel), and $(2)$ estimate the parameters of these two sets of functions [^2] In a fully Bayesian setting, instead of estimating the parameters, their posterior distributions are obtained and pblackictions on $\boldsymbol{y}_{q}$ require marginalization with respect to these distributions. Due to the significantly higher computational costs of fully Bayesian techniques and their marginal accuracy improvements in the case of GPs, we recommend and use MAP..
More formally, assume the training dataset $\mathopen{}\left\{\boldsymbol{x}^{(i)},y^{(i)}\right\}\mathclose{}_{i=1}^{n}$ is given where $\boldsymbol{x}=[x_{1},...,x_{dx}]^{T}\in\mathbb{X}\subset\mathbb{R}^{dx}$ and $y^{(i)}=y(\boldsymbol{x}^{(i)})\in\mathbb{R}$ denote the inputs and response [^3] We focus on regression problems whose output dimensionality is one but note that GPs can handle multi-response or multi-task problems as well [RN1902, RN328, RN287]., respectively. Given $\boldsymbol{y}=[y^{(1)},\cdots,y^{(n)}]^{T}$ and $\boldsymbol{X}$ whose $i^{th}$ row is $(\boldsymbol{x}^{(i)})^{T}$, our goal is to pblackict $y(\boldsymbol{x}^{*})$ at the arbitrary point $\boldsymbol{x}^{*}\in\mathbb{X}$. Following the above description, we assume $\boldsymbol{y}=[y^{(1)},\cdots,y^{(n)}]^{T}$ is a realization of a GP with the following parametric mean and covariance functions:
@@
where $\omega_{i}\in\color[rgb]{0,0,0}{\mathbb{R}}$ [^4] To ensure numerical stability, $\omega_{i}$ is typically bounded to a subset of $\color[rgb]{0,0,0}{\mathbb{R}}$, e.g., $[-10,4]$., $p\in[1,2]$, $\nu\in\{\frac{1}{2},\frac{3}{2},\frac{5}{2}\}$, $K_{\nu}$ is the modified Bessel function of the second kind, and $\Gamma$ is the gamma function. The inductive bias that the kernels in Equation 4c encode into the learning process is that close-by input vectors $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$ have similar (i.e., correlated) output values. The degree of this correlation is quantified by the interpretable length-scale (aka roughness) parameters where the magnitude of $10^{\omega_{i}}$ is directly related to the response fluctuations along $x_{i}$.
@@
where $c(\boldsymbol{x}^{*},\boldsymbol{X};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})$ is a $1\times n$ row vector with entries $c_{i}=c(\boldsymbol{x}^{*},\boldsymbol{x}^{(i)};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})$ and its transpose is $c(\boldsymbol{X},\boldsymbol{x}^{*};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})$. Equations 9a and 9b can be straightforwardly extended to pblackict the response distribution for a batch of samples denoted by $\boldsymbol{X}^{*}$:
@@
$\text{cov}{\color[rgb]{0,0,0}\mathopen{}\left(y\mathopen{}\left(\boldsymbol{X}^{*}\right)\mathclose{},y(\boldsymbol{X}^{*})\right)\mathclose{}}=c(\boldsymbol{X}^{*},\boldsymbol{X}^{*};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})-c(\boldsymbol{X}^{*},\boldsymbol{X};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})\boldsymbol{C}^{-1}c(\boldsymbol{X},\boldsymbol{X}^{*};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})$ (10b)
The above formulations build interpolating GPs. To handle datasets with noisy observations, the nugget or jitter parameter, denoted by $\delta$ [RN1917, RN332, RN1908], is used where $\boldsymbol{C}$ is replaced by $\boldsymbol{C}_{\delta}=\boldsymbol{C}+\delta\boldsymbol{I}_{nn}$ where $\boldsymbol{I}_{n\texttimes n}$ is the $n\times n$ identity matrix (with this adjustment, the stationary noise variance estimated by the GP is $\widehat{\delta}$) [^6] Some recent works [RN292, bostanabad2018leveraging, macdonald2015gpfit] apply the nugget directly to $\boldsymbol{R}$ but herein we adhere to [RN1917] and add $\delta$ to $\boldsymbol{C}$.. In addition to modeling stationary noise, the nugget parameter is also used to mitigate the numerical issues associated with $\boldsymbol{C}$. That is, even with noise-free $\boldsymbol{y}$, $\boldsymbol{C}_{\delta}$ is used while minimizing Equation 8 to ensure the correlation matrix is always invertible. When the nugget parameter is used for fitting a GP to noisy observations, Equation 1 takes on the following form:
@@
$\text{cov}\mathopen{}\left(y\mathopen{}\left(\boldsymbol{X}^{*}\right)\mathclose{},y(\boldsymbol{X}^{*})\right)\mathclose{}=c(\boldsymbol{X}^{*},\boldsymbol{X}^{*};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})-c(\boldsymbol{X}^{*},\boldsymbol{X};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})\boldsymbol{C}^{-1}_{\delta}c(\boldsymbol{X},\boldsymbol{X}^{*};\widehat{\boldsymbol{\theta}},\widehat{\sigma}^{2})+\widehat{\delta}\boldsymbol{I}.$ (12b)
We highlight that Equation 12b does not consider the additional uncertainties incurblack by estimating the parameters of the mean and covariance functions (note that Equation 2 assumed the mean vector and covariance matrices are known). These additional uncertainties can be quantified by building and using the GP within a Bayesian framework where sampling methods (e.g., Markov Chain Monte Carlo or MCMC) are requiblack for marginalization as closed-form expressions are only available for specific cases (see [RN1896] for an example). Since such sampling methods are typically expensive and the provided benefits are marginal, MAP is frequently employed in GP modeling [murphy2012machine].
@@ 3 Related Works
Many open-source GP libraries have been recently developed and in this section we review some of the most well-known ones. One of the earliest open-source GP packages is TreedGP [gramacy2007tgp] which is developed in R and primarily aims to address the stationarity and scalability issues of GPs. In particular, TreedGP recursively partitions the input space via parallel and axes-aligned boundaries [chipman1998bayesian] and then endows each partition with a GP whose covariance function is stationary. TreedGP uses Bayesian averaging to combine these GPs which is particularly important for obtaining smooth predictions on the partition boundaries. The major limitations of TreedGP are its non-differentiability on the boundaries, high computational costs (as the Bayesian analyses rely on MCMC [harkonen2022mixtures]), reliance on trees which can only partition the input space with axis-aligned boundaries [candelieri2021treed], and inability to efficiently handle categorical features in small-data applications.
GPfit [macdonald2015gpfit] and GPM [bostanabad2018leveraging] are also R packages and they are primarily designed to improve the hyper-parameter optimization process at the expense of increased computational costs. GPfit has a multi-step pre-processing stage that aims to improve the quality of the initial points that are used via L-BFGS [^7] Limited-memory Broyden–Fletcher–Goldfarb–Shanno in minimizing Equation 8. Unlike GPfit, GPM develops a multi-step continuation-based strategy to increase both the robustness and accuracy of the optimization process. In particular, GPM indirectly controls $\delta$ via the auxiliary parameter $\epsilon$ that puts a lower bound on the smallest eigenvalue of the correlation matrix [^8] The rationale behind this choice is that the smallest eigenvalue of $\boldsymbol{R}$ can sometimes be negative due to numerical issues.. GPM first uses a large value for $\epsilon$ (e.g., $10^{-2}$) and minimizes Equation 8 while requiring the smallest eigenvalue of $\boldsymbol{R}$ to always be larger than the imposed $\epsilon$. In addition to guaranteeing numerical robustness, this requirement dramatically smooths the profile of the objective function and hence most (if not all) optimizations quickly converge to the same solution. Then, GPM relaxes the constraint on $\boldsymbol{R}$ (e.g., $\epsilon=10^{-3}$) and repeats the optimization while using the solution(s) of the previous step as the initial guess(es) in the current step, see Figure 1. These steps are continued until the minimum value of $\epsilon$ is reached and then the parameters of the final GP are chosen by identifying the step (or $\epsilon$) at which the leave-one-out cross-validation (LOO-CV) error of the model is minimized.
@@
[…]
Other notable open-source GP libraries that we mention in passing include those of Ambikasaran et al. [RN1898] which hierarchically factors the covariance matrix [^9] As long as it is built with specific covariance functions such as the Gaussian or Matérn. into a product of block low-rank updates of the identity matrix to accelerate matrix inversion and determinant calculation (their method loses accuracy for $dx>3$), Vanhatalo et al. [RN1900] which is a Matlab library that integrates various elementary computational tools (e.g., sparse approximation) for GP-based regression, and GPML [RN332] which is also a Matlab library and has been widely used in a wide range of applications.